Kanata Construction specializes in large projects in Edmonton and Saskatoon. In 2017, Kanata invested $1.6 million in new excavating equipment, which qualifies for a CCA rate of 50%. At the same time the firm sold some older equipment on the secondhand market for $210,000. When it was purchased in 2014, the older equipment cost $600,000.

Calculate the UCC for the asset pool in each year from 2014 through 2018. (Round the final answers to 2 decimal places. Omit $ sign in your response.)

Year CCA Ending UCC
2014 $ $
2015 $ $
2016 $ $
2017 $ $
2018 $ $
Answer #1
Year CCA $ CCA working Ending UCC $ Ending UCC Working
2014 150,000.00 =600000*50%*1/2    450,000.00 =600000-150000
2015 225,000.00 ( 450,000 x 50% )    225,000.00 (450,000 - 225,000 )
2016 112,500.00 (225,000 x 50% )    112,500.00 (225,000 - 112,500 )
2017 400,000.00 (1,600,000 x 50% x 1/2 ) 1,200,000.00 =1600000-400000
2018 600,000.00 (1,200,000 x 50% )    600,000.00 (1,200,000 - 600,000 )